OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

Plug g(x) in for x

OpenStudy (jack1):

f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). f(g(x)) = 9(-x + 3) - 2

OpenStudy (anonymous):

how did you get that?

OpenStudy (jack1):

f(g(x)) = 9(-x + 3) - 2 = -9x + 27 - 2 how we got it was the question itself it was saying if z = 9y - 2 ... and y = 3 - x .... what does z equal in terms of x...?

OpenStudy (anonymous):

ohh okay now I understand thank you can you help me with another one but this time ill try and do it and you see if im write or wrong :)

OpenStudy (anonymous):

Compositions of functions simply take one function (in it's entirety) and substitute it in as the variable in the initial function. F(x)=9x-2... x is the variable. g(x)= -x+3 is the function that will sub for the variable. So F(g(x))= 9(-x+3)-2

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

you would write it like -4x(10x-6) +7 right?

OpenStudy (jack1):

yep

OpenStudy (anonymous):

f(g(x))=-4(g(x))+7

OpenStudy (jack1):

no remove the x

OpenStudy (jack1):

-4(10x-6) +7

OpenStudy (anonymous):

How would you remove the x

OpenStudy (anonymous):

what happened to the x, how did you cancel it out

OpenStudy (jack1):

x of first equation is )10x-6)

OpenStudy (anonymous):

oh okay so it would be -4(10x - 6) +7 -40x + 24 + 7 -40x + 31

OpenStudy (jack1):

think of it this way: if z = -4y + 7 ... and y = 10x - 6 .... what does z equal in terms of x

OpenStudy (jack1):

but yeah, perfect ^^^

OpenStudy (anonymous):

yeah! are you willing to help with another one

OpenStudy (jack1):

ok one mores

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (jack1):

go fur it dude...

OpenStudy (anonymous):

I'm a little stuck what does it mean by f(g(2))?

OpenStudy (jack1):

* sorry if z = y + 8 ... and y = x^2 - 6x - 7 .... what does z equal in terms of x, when x = 2 so first do z in terms of x

OpenStudy (anonymous):

f(x)= x+8 f(x)= 2 + 8 f(x) = 10 g(x) = x2 - 6x - 7 g(x) = 4-12-7 g(x) = -15

OpenStudy (jack1):

sorry dude... f(x) = y + 8 and g(x) = y so y = x^2 - 6x - 7 so when x = 2... y =4-12-7 y = -15 so u got the 2nd half right...

OpenStudy (jack1):

now sub y = -15 into f(x) = y + 8

OpenStudy (anonymous):

-15 + 8 -7

OpenStudy (jack1):

yep, perfect it was my bad bru, i mis-typed in the earlier post

OpenStudy (anonymous):

its okay thanks for all your help!!!

OpenStudy (jack1):

ALL GOOD DUDE, SLATERS